OpenStudy (anonymous):

A city population, which was initially 15,500, has been dropping by 3% a year. Write an exponential function and graph the function. Use your graph to predict when the population will drop below 8,000. Could someone help me solve this?

OpenStudy (whpalmer4):

If it is dropping by 3% each year, that means after 1 year, it has 97% of the previous year's population, or 0.97*previous year's population. If we take \(P_0\) to be the initial population, after 1 year, we have \(P(1) = P_0*0.97\) After two years, we have \(P(2) = P(1)*0.97 = P_0*0.97*0.97 = P_0(0.97)^2\) After three years, we have \(P(3) = P(2)*0.97 = P_0(0.97)^2*0.97 = P_0(0.97)^3\) Do you see a pattern developing here?

OpenStudy (anonymous):

yess

OpenStudy (whpalmer4):

Okay, so what is your formula for \[P(t)=\]

OpenStudy (anonymous):

P(t) = 15,500(0.97)t Right??

OpenStudy (anonymous):

So then, the population will drop below 8,000 in 21.7 years right?

OpenStudy (whpalmer4):

That appears to be correct!

OpenStudy (anonymous):

Awesome! That helps a lot! Thank you

OpenStudy (anonymous):

:)

OpenStudy (the_fizicx99):

I thought about graphing it myself in my head being a simple function eh, but I forgot what the decay rate would be like >_< doing it algebraically would also give me the same answer but graphing it is faster.

OpenStudy (the_fizicx99):

Nvm rolf just answered myself ._.

OpenStudy (anonymous):

hahaha :) silly